protons in uniform magnetic field of .3T follow a circular trajectory with 20cm radius. Determine the speed of the protons. (q=1.6×10 -¹⁹ C and m=1.673×10 -²⁷ kg)

1
Expert's answer
2022-04-05T09:44:28-0400

The radius of the circular path is

"r=\frac{mv}{Bq}"

The speed of the proton is

"v=\frac{Bqr}{m}"

"=\frac{(0.3)(1.60\times10^{-19})(0.20)}{1.673\times10^{-27}}"

"=5.738\times 10^6 m/s"